OpenStudy (anonymous):

A bus moves away from rest at a bus stop with an acceleration of 1ms^-2. As the bus stars to move a man who is 4m behind the stop runs with a constant speed after the bus. If he just manages to catch the bus find his speed?

OpenStudy (anonymous):

hey.., what's the value of speed of a bus??

OpenStudy (anonymous):

That is all the question

OpenStudy (anonymous):

ok..., s = vo.t + 0.5 a t^2 4 = 0 + 0.5 (1). t^2 t^2 = 4/0.5 t^2 = 8 t = sqrt(8) vt = vo + at vt = 0 + (1 x sqrt(8))

OpenStudy (anonymous):

Yeah that fits the answer which is root8, thanks a lot

OpenStudy (anonymous):

welcome